OpenStudy (anonymous):

y-6=1/4(x+8) solve for y

OpenStudy (anonymous):

\[y=1/4x+8\]

OpenStudy (anonymous):

is this the correct answer

OpenStudy (anonymous):

Just in slope intercept form, yes? y-6 = 1/4(x+8) y - 6 = 1/4x +2 y - 6 +6= 1/4x +2 +6 y = 1/4x +8

OpenStudy (mathmale):

Looks fine to me!

OpenStudy (anonymous):

oh thank you so much